OpenStudy (anonymous):

a reaction produces 0.883 moles of H2O. How many molecules are produced?

OpenStudy (anonymous):

Remember this grams-moles-atoms

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh okay remember that got it!

OpenStudy (anonymous):

so advogardos constant 6.022*10^23 for 1atom

OpenStudy (anonymous):

i know that part

OpenStudy (anonymous):

so 0.883molesH2O* 6.022*10^23=5.32*10^23

OpenStudy (anonymous):

so is 5.32 x 10^23 the answer?

OpenStudy (anonymous):

we cancel the bottom .883molesH20*6.022^23= answer 1mole H20

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Remember Gram-Moles-Atoms

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

np did you understand

OpenStudy (anonymous):

yes